40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 537     Entry time: Wed Jun 18 00:19:29 2008
Author: rob 
Type: Update 
Category: PSL 
Subject: MOPA trend 
15 day trend of MOPA channels. The NPRO temperature fluctations are real, and causing the PMC to consistently run up against its rails. The cause of the temperature fluctations is unknown. This, combined with the MZ glitches and Miller kicking off DC power supplies is making locking rather tetchy tonight. Hopefully Yoichi will find the problem with the laser and fix it by tomorrow night.
Attachment 1: MOPAtrend.png  119 kB  | Show | Show all
ELOG V3.1.3-